OpenStudy (anonymous):

the sum of two numbers is 52.their difference is 36 find the two numbers. help!

OpenStudy (tiffany_rhodes):

Let x+y=52 and x-y=36. Use system of equations to cancel out the y's. So 2x=88 x=88/2 which equals 44. So x=44. Plus x back into to either equation and solve for y. I plugged x=44 into the equation x+y=52 and found y=8. Hope that helps! Double check my math.

OpenStudy (anonymous):

Thank you!

OpenStudy (tiffany_rhodes):

yeah no problem!

OpenStudy (anonymous):

(: